### v1

# Generate E-commerce Product Background by Integrating Category Commonality and Personalized Style

Abstract
The state-of-the-art methods for e-commerce product background generation suffer from the inefficiency of designing product-wise prompts when scaling up the production, as well as the ineffectiveness of describing fine-grained styles when customizing personalized backgrounds for some specific brands. To address these obstacles, we integrate the category commonality and personalized style into diffusion models. Concretely, we propose a Category-Wise Generator to enable large-scale background generation for the first time. A unique identifier in the prompt is assigned to each category, whose attention is located on the background by a mask-guided cross attention layer to learn the category-wise style. Furthermore, for products with specific and fine-grained requirements in layout, elements, etc, a Personality-Wise Generator is devised to learn such personalized style directly from a reference image to resolve textual ambiguities, and is trained in a self-supervised manner for more efficient training data usage. To advance research in this field, the first large-scale e-commerce product background generation dataset BG60k is constructed, which covers more than 60k product images from over 2k categories. Experiments demonstrate that our method could generate high-quality backgrounds for different categories, and maintain the personalized background style of reference images. The link to BG60k and codes will be available soon.
1 Introduction
E-commerce product background generation is practically valuable in the research of image generation. It aims to generate natural and realistic backgrounds for specific products, thereby enhancing the online performance like click-through rate (CTR) of advertising images [3, 30, 13]. Due to the significant differences in the background of advertising images from different categories and brands, advertisers traditionally employ professional designers to design suitable backgrounds for their products, which is often time-consuming and financially burdensome.
Figure 1: (a) Large-scale advertising images from the e-commerce platform. (b) Advertising images clustered by the category. Images in the same categories share the similar background. (c) Advertising images clustered by the brand. In addition to the category-wise common style, images of some brands have their signature styles.
To reduce the cost of advertisers, most previous methods [13, 4, 5, 24] pre-generate a complete image, which is then combined with the product and other visual elements to obtain an advertising image. As the background is independent of the appearance of the product during generation, the final image exhibits a discernibly low level of realism. To address this issue, recently text-based inpainting methods [21, 16, 1, 2, 29, 14, 15, 23] are widely applied, where the background is generated with the consideration of the product. However, it is still time-consuming to design a proper prompt for each product to describe the content of the background, harming the efficiency in large-scale background generation, as shown in Fig. 1 (a). Besides, products of some specific brands demand backgrounds in a fine-grained and consistent personalized style, which is almost ineffective to describe only by text.
For achieving large-scale background generation efficiently, we integrate the category commonality into diffusion models. We observe that the intra-category variance among the background styles is significantly less than the inter-category variance. As shown in Fig. 1 (b), the backgrounds of laptops usually protrude a sense of modern technology, while the rice cookers are mainly on the dining tables. Therefore it is natural to consider simplifying the prompt design by leveraging such category commonality. Concretely, we present a Category-Wise Generator, where we templatize the prompt and assign an identifier to each category for its common background style. A mask-guided cross attention layer is also devised to enforce the attention of identifier only on the background. The identifier could finally serve as the alternative to the prompt designed for the specific category during background generation.
Nevertheless, beyond the common character in the category, the advertising images of some well-known brands often exhibit their fine-grained signature styles in the visual effect formed by the comprehensive influence of the layout, elements, color, lighting, etc. As shown in Fig. 1 (c), in addition to the sense of technology, the background of Alienware is mainly purple, while Omen is mainly red. Therefore, we further propose a Personality-Wise Generator in parallel with the Category-Wise Generator, to incorporate the personalized style into the category-wise general background. Due to the poor effectiveness of textual representation in describing fine-grained or abstract information, the Personality-Wise Generator generates personalized backgrounds by directly mimicking the characters of the reference advertising images. To leverage the training data more efficiently, we adopt the self-supervised training to consider an advertising image as both the reference and ground truth image. We also perturb the reference image by a series of augmentation strategies during training, which is crucial to prevent the generated background from a simple copy of the reference image during inference.
To the best of our knowledge, there is no publicly available dataset for training large-scale product background generation methods. Therefore, we elaborately mine and clean a large amount of advertising images from the famous E-commerce platform to fill in this gap. The finally released dataset covers over 2k categories, and contains more than 60k images with their category names and product masks.
Our contributions are summarized as follows:
  • We present a Category-Wise Generator to enable the efficient large-scale background generation for the first time. Specifically, we assign a unique identifier to each category and design a mask-guided cross attention layer to locate the attention of the identifier only on the background. The identifier could finally serve as the prompt to generate background in the common style within each category.
  • We further propose a Personality-Wise Generator to maintain the personalized styles of some brands by mimicking the characters of reference images. Considering the lack of similar background pairs during training, we employ self-supervised training as well as a background perturbation strategy.
  • We release the first large-scale dataset namely BG60k to advance the research of the product background generation. Extensive experiments demonstrate that our method outperforms the state-of-the-art inpainting methods by a large margin in both background similarity and quality.
2 Related Work
2.1 Background Generation
Given the diversity of e-commerce product backgrounds, advertisers often need to employ designers to produce attractive images. Suffering from the consequent expensive cost, several methods [13, 4, 5, 24] combine various elements based on templates, where the background is an entire image generated in advance. Mishra et al. [13] explore the possibility of collaboratively learning advertisement creative refinement via A/B tests of multiple advertisers. AES [4] utilizes dynamic programming and Thompson sampling to efficiently select visual elements with the highest CTR. AutoCO [5] applies factorization machine (FM) and one-shot search algorithms to model complex interactions between visual elements. However, these template-based methods do not consider the appearance of the product when generating the background, resulting in a discernible disconnection between foreground and background elements. Recently, Ku et al. [10] leverage pix2pix [9] to generate an image background for a product, where the background is customized based on the appearance of the product. Unfortunately, it only targets one category and is not suitable for large-scale e-commerce scenarios. In contrast, our method covers thousands of categories and is thus more practically valuable.
2.2 Image Inpainting
According to the requirements of the advertising images, the appearance of the product needs to remain strictly unchanged. Therefore, only the background area needs to be generated, which is similar to the image inpainting task [27, 8, 28, 21, 12]. To increase the receptive field, dilated convolutions [8, 27] are used to adapt to large and irregular hole inpainting. Contextual attention [28] explicitly leverages surrounding image features as references during training. LaMa [21] employs Fourier Convolutions to integrate both global and local information efficiently. With the rise of diffusion models in the field of image generation, RePaint [12] applies Denoising Diffusion Probabilistic Models (DDPM) [7] to image inpainting, which alters the reverse diffusion iterations by sampling the unmasked regions using the given image information.
Figure 2: The overall framework of our methods during training. CG, SD and PG are short for the Category-Wise Generator, pretrained Stable Diffusion and Personality-Wise Generator, respectively. During inference, the perturbation is removed. Notably, without specific requirement on the fine-grained personalized style, CG could work solely without PG.
A different line of research is guiding image inpainting with texts [16, 1, 2, 29, 14, 15, 23] or images [26, 25, 11, 22]. Stable Diffusion Inpainting [16] is a text-to-image model that generates realistic images based on text input, utilizing a UNet [17] architecture with additional input channels for the masked-image and mask. Blended Diffusion [1] performs local edits using natural language descriptions and RoI masks. ControlNet Inpainting [29] controls pretrained Stable Diffusion conditioned on inpainted images and texts. However, these methods require laborious prompt engineering for detailed description. Yang et al. [26] argue that images can better express concrete ideas, and paint with a conditional image. Nevertheless, the images used to drive this method are inherently related to the foreground, whereas our work specifically focuses on the background. The background, being typically more abstract and containing multiple objects, poses a greater challenge in generation.
3 Method
We generate the background by inpainting to ensure the authenticity of the product and the overall harmony of the image. The overall framework is illustrated in Fig. 2, which consists of three components: Stable Diffusion (SD), the Category-Wise Generator (CG) and the Personality-Wise Generator (PG), where the latter two modules are built and modified based on the architectures of ControlNet [29] (3.1). During training, given the advertising image $I$ and the mask of the product $M$, CG takes as the input $I\otimes M$ to integrate the general category knowledge (3.2), while PG captures the personalized style from $I\otimes(1-M)$ (3.3), where $\otimes$ is Hadamard product. Then the multi-level outputs of the two generators are all incorporated into the SD Decoder in a similar way to ControlNet to predict the latent noise.
3.1 Preliminaries
We follow the paradigm of ControlNet to design our whole pipeline for the sake of its hierarchic and sufficient control of image generation. ControlNet works as a parallel branch of SD to guide such text-to-image model to adapt for various image-to-image tasks through an extra conditional image. It is composed of an SD Encoder and several zero convolutions, and shares the same encoded prompt and latent noise with SD. Specifically, given the input conditional image, the feature output from each block in the ControlNet encoder is added to the feature from the positionally symmetrical SD decoder block. During training, ControlNet is optimized by minimizing the following objective:
$\mathbb{E}_{z,c,\epsilon\sim\mathcal{N}(0,1),t}\Bigl{[}\|\epsilon-\epsilon_{\theta}(z_{t},t,P,c)\|^{2}_{2}\Bigr{]},$ (1)
where $P$, $c$, $z_{t}$ and $\theta$ are the encoded prompt, the conditional image, the latent noise at the time step $t$, and the trainable parameters of ControlNet, respectively. During inference, the random latent noise $z_{T}$ sampled from $\mathcal{N}(0,1)$ is denoised iteratively to $z_{0}$ based on the encoded prompt and conditional image. Then $z_{0}$ is further decoded to the final generated image by the pretrained VAE.
3.2 Generation with Category Commonality
In this subsection, we focus on the efficient large-scale background generation. Although the e-commerce platforms cover diverse products, we notice that the backgrounds of the products in the same category have many characters in common. Therefore it is natural to consider sharing the same prompt among the products in the same categories, e.g., injecting the category name into a prompt template ’A photo of [category]’. This reduces the complexity of prompt engineering to $\mathcal{O}(1)$. Then we could design CG by adopting the same architecture as ControlNet, and taking as input the prompt template and the product image with masked background. After trained on images clustered by the categories, i.e., BG60k, CG is able to summarize the category commonality. In this way, the category name in the prompt is mapped to the category commonality for generating the background in the corresponding style.
However, it is sub-optimal to simply store the category knowledge into the category name in prompt, as well as inherit the original architecture of ControlNet. The category name should have been utilized to generate backgrounds, but the knowledge of the foreground is also encoded into it unavoidably during the foreground generation in training. We handle this problem from two aspects. Firstly we divide the prompt into two parts: the product prompt describing the product, and the background prompt describing the background. Concretely, inspired by Dreambooth [18], we map each category name into an identifier and construct the prompt as ’A photo of [category], in the background of [D]’, where [D] is the specific identifier. Secondly we further devise a mask-guided cross attention layer to ensure that the product and background generation solely rely on their corresponding prompt respectively. Specifically, the original cross attention layer could be formulated as:
$x_{out}^{i}=x_{in}^{i}+CA(x_{in}^{i},P),$ (2)
where $CA(\cdot)$, $x_{in}^{i}$ and $x_{out}^{i}$ are the cross attention layer, the input and output of the $i_{th}$ cross attention layer, respectively. Then the mask-guided cross attention layer is formulated as:
$x_{out}^{i}=x_{in}^{i}+CA(x_{in}^{i},P_{fg})\otimes M+CA(x_{in}^{i},P_{bg})\otimes(1-M),$ (3)
where $M$, $P_{fg}$, $P_{bg}$ indicate the mask of the product, the encoded product prompt (’A photo of [category]’) and the background prompt (’in the background of [D]’), respectively. We replace all the attention layers in CG with the mask-guided ones, forcing the sub-network to generate the background based on the background prompt only. Similar to ControlNet, the $x_{out}^{i}$ from all cross attention layers is finally added to the feature output from the positionally symmetrical SD decoder block, to integrate the category commonality and the product feature into SD Decoder. During inference, [D] could serve as the prompt to describe the common background style of a specific category.
3.3 Generation with Personalized Style
The backgrounds containing category-wise commonality are appropriate for most products, but not for some well-known brands, which require their consistent signature background styles. Hence superposed on the category-wise background, we further propose PG to generate the personalized background similar to the reference image in the layout, elements, etc. Practically the reference images come from the existing advertising images within the same brand (or even from other brands if possible). Nevertheless, it is non-trivial to capture such information from the reference image, as it usually contains multiple elements, more complex spatial relationships, and another undesired product. Hence we also adopt the similar architecture of ControlNet, as the output high resolution feature maps (rather than the single token [26]) preserve the semantic cues with spacial information naturally. Moreover, to confine the personalized style only on the generated background, we mask the output $y_{out}^{i}$ of the $i_{th}$ cross attention layer by:
$y_{out}^{i}:=y_{out}^{i}\otimes(1-M).$ (4)
Due to the unified diffusion structure, PG and CG share the same number of encoder blocks. Hence we add each $y_{out}^{i}$ to $x_{out}^{i}$ when integrated into SD Decoder, to generate backgrounds with incorporated personalized style.
As there lack enough training pairs comprising a reference image and a similar ground-truth generated image, PG is trained in a self-supervised fashion. During training, we sample an advertising image in each iteration and obtain its product and background. The product and the background are fed to CG and PG to reconstruct the image itself. In this way, the ground truth of the generated background is the original background itself. Such a simple mechanism leads to a degradation in the model’s functionality, since the PG could find a shortcut and learn to paste the product directly onto the background. In practical applications, the product in the reference image often exhibits a different shape from the one requiring a background. Consequently, employing a copy-and-paste approach may result in disharmony within the generated image.
Therefore, we resort to a background perturbation strategy against the copy-and-paste. Specifically, supposing the original advertising image and the mask of the product are denoted as $I$ and $M$, we firstly perturb $M$ by:
$M_{aug}=Dilate(M|\theta),$ (5)
where $Dilate(\cdot|\theta)$ is the dilation function with kernel $\theta$. In this way, PG learns to tackle the difference between the original product and the product in the reference image by inpainting the region $I\otimes(M_{aug}-M)$. For a similar reason, we apply other augmentation strategies like adding random masks and mask translation. Furthermore, we leverage the mixup strategy to overlay a random advertising image onto the reference image, in order to penalize the PG for simply replicating the elements from the reference image. Specifically, $I$ is perturbed by:
$I_{aug}=\sigma\otimes I+(1-\sigma)\otimes I_{rand},$ (6)
where $I_{rand}$ is another randomly sampled advertising image and $\sigma$ is the hyper-parameter. We feed $I_{aug}\otimes M_{aug}$ to PG instead. It turns out that PG learns to understand more about the elements in the reference image, filtering out the inappropriate portion and generating semantically similar elements for the other portion.
Although the background perturbation strategy prevents PG from the copy-and-paste, we experimentally find that it has a negative impact on the quality of the generated images. To resolve this obstacle, we utilize the original high-quality reference image as the initial latent image during inference, i.e., replacing $z_{T}$ with the latent reference image with noise $\epsilon\sim\mathcal{N}(0,1)$ added for $T$ times. This reduces the difficulty of generating images similar to reference images, and is thereby beneficial for improving the image quality.
4 BG60k
4.1 Motivation
Figure 3: Some unsatisfying examples in LAION-5B [19]. Models trained with such images will generate images in poor quality.
It is commonly known that the performance of the generation models relies heavily on the quality of the training data. Recently, LAION-5B [19] and its several subsets might be the most widely-used publicly available dataset for the training of diffusion models such as Stable Diffusion. Despite the significant scale of the training images, there still exist gaps between LAION-5B and the advertising images on the e-commerce platforms.
Concretely, LAION-5B is not designed specifically for e-commerce scenarios, and thus many training images do not meet the requirements of advertising images. Taking the product washing machine as an example, we retrieve some images with high aesthetics scores in LAION-5B by CLIP [^1] https://laion.ai/blog/h14_clip_retrieval/ in Fig. 3. It could be found that the washing machines in these images might be too small (Fig. 3 (a)), placed in the non-salient area (Fig. 3 (b)), or viewed at an improper perspective (Fig. 3 (c)). Meanwhile the background could be dusky (Fig. 3 (b)), messy (Fig. 3 (e)) or contain many worn-out objects (Fig. 3 (d)). Models trained on such dirty data will fail to generate satisfying advertising images, as illustrated in Fig. 5 (b). Although other work [10] collects images in an advertising platform instead, the dataset only covers one category and is not publicly available.
4.2 Dataset Construction and Analysis
In order to reduce the above gaps, we construct and release the first large-scale e-commerce background generation dataset BG60k. We first mine millions of advertising images from the famous e-commerce platform. Then we clean the data based on the following requirements:
  • Appealing: To ensure the quality of the dataset, we only retain images with high click-through rate and not in blank background.
  • No texts: Texts and logos in images are annotated and erased by LaMa [20], as they are not our focus. We then review the inpainted images to ensure their harmony.
  • No body: Considering the right of portrait, we delete images containing human bodies or faces.
Finally we obtain 63293 advertising images from 2032 categories and associate each image with its corresponding category. Therefore each item in the dataset is composed of an original advertising image $I$, a mask of the product $M$ and the category of the product.
Moreover, we build two extra testing datasets to evaluate our methods comprehensively. Concretely, we choose 1k products and their original backgrounds from over 200 categories, namely BG1k. This allows us to compare the difference between the generated and original background statistically. Additionally, we select 1.6k pairs of product and reference image to evaluate the ability of generating with personalized style, namely BG-pair. Each image pair is composed of a product image to generate the background, the category name, the reference image and the mask of the original product in it. We check each image pair to ensure the feasibility of generating a similar background.
Figure 4: Examples in BG60k. Compared with images in Fig. 3, the training images in BG60k are more appealing and exquisite.
The image number of the categories in BG60K follows the long-tail challenging distribution of the real advertising images with high impressions, since there are around $6\%$ of categories containing over 100 images, while over $50\%$ of categories less than 10 images. The detailed distribution of categories in BG60k could be found in the supplementary materials. We also visualize some examples of BG60K in Fig. 4, which proves that the training images are exquisite and exhibit prominent products.
5 Experiments
5.1 Implementation Details
Our model is trained on BG60k for 300 epochs with a learning rate 1e-5. The total batch size is 128 and each image is resized to $512\times 512$. The identifier is implemented with the combination of sks [18] and the id of the category. For data augmentation, the size of $\theta$ is 50, and the $\sigma$ in mixup is sampled from a uniform distribution $\mathcal{U}(0.75,0.95)$. Other hyper-parameters of the training are consistent with ControlNet. During inference, we adopt a 50-step DDIM sampler with the classifier-free guidance scale 2.0.
5.2 Evaluation Metrics
Quantitative Metrics: The generated backgrounds are evaluated from two aspects. On the one hand, the background should be exquisite, either photorealistic (for products like home appliances) or artistically stylized (for products like cosmetics). It should also be harmonious with the product to make the entire image look like an advertising image. Following Yang et al. [26], we measure this aspect by FID score [6], which reflects the distribution gap between features of two groups of images. On the other hand, the background of the generated image is supposed to be similar to the reference image (if given). This is quantified by the similarity of CLIP visual features of the two images.
User Study: We further conduct the user study for a more comprehensive comparison. Concretely, we sample one hundred product-reference pairs from BG-pair, and combine the original product, the reference image, and images generated by our method and counterparts. The order of the generated images is shuffled randomly for each group. Then we invite twenty experienced advertising practitioners to rank the consistency with the reference image and the quality of the generated background. When evaluating the background consistency, the practitioners are mainly concerned about the layout, elements, color and overall style. When evaluating the image quality, the practitioners are particularly concerned about the overall aesthetics, the harmony and the suitability of serving as the advertising image.
5.3 Comparison with State-of-the-Art
5.3.1 Large-Scale Background Generation
In this subsection, we evaluate our method for efficient large-scale background generation. Without extra condition input (the reference image) included, we remove PG and compare CG with the state-of-the-art inpainting methods Stable Diffusion and ControlNet, as well as an unconditional inpainting method LaMa [20] on BG1k. Since it is not feasible to carefully devise the prompt for each product, the backgrounds are generated based on the prompt template. Specifically, we adopt ’A photo of [category], in the background of [D]’ as the prompt template for CG, and ’A photo of [category]’ for Stable Diffusion and ControlNet, as [D] is the garbled character for them.
Table 1 summarizes the comparison on CLIP similarity (the generated image v.s. the original image) and FID. It could be found that CG has the superiority of 1.25 CLIP similarity and 1.85 FID. It means that the backgrounds generated by CG are more like the original high-quality backgrounds and thus better meet the category-wise requirements of the advertisements in the large-scale generation.
Method CLIP Similarity ($\uparrow$) FID ($\downarrow$)
LaMa [20] 88.33 8.24
Stable Diffusion [16] 76.47 7.54
ControlNet [29] 84.80 5.79
Ours 89.58 3.94
Table 1: Comparison with image inpainting methods conditioned on the prompt template on BG1k.
Figure 5: (a) tSNE visualization, where the circle represents the cluster center in training data. Each triangle/square indicates the background embedding of one image generated by CG/ControlNet. (b) Visualization of comparison on the category refrigerator. Compared with ControlNet, CG could generate more realistic and exquisite backgrounds.
For a more comprehensive analysis on whether our generated backgrounds contain the category commonality, we visualize the features of backgrounds from our method and ControlNet by tSNE in Fig. 5 (a). It could be seen that the features of backgrounds generated by CG are clustered around the corresponding centers more compactly. On the contrary, background features generated by ControlNet are farther away from the centers, and even contain some outliers. Furthermore taking the category ’refrigerator’ as an example, we compare the generated images in Fig. 5 (b). It turns out that ControlNet might generate backgrounds lacking photorealism (the first column), irrelevant to the product (the third column), or containing other product of the same category (the fifth column), while CG is able to generate indoor backgrounds consistently. This indicates that CG is able to summarize the rule that the refrigerator is better placed in the room in the advertising image.
5.3.2 Personalized Background Generation
In this subsection, we consider the fine-grained personalized background generation for some specific brands conditioned on the reference advertising image. Specifically, for each pair of product and reference image from BG-pair, we compare our generated backgrounds with those from the following methods: 1) Blended Diffusion [1]: we leverage the loss between the background features of reference and generated image as the classifier-guidance loss. 2) Paint-by-Example [26]: we encode the reference image as the condition. 3) Stable ${}^{\dagger}$ [16]: we first paste the product on the reference image, and then initialize $z_{T}$ with the noisy latent adjusted reference image. 4) ${}^{\dagger}$ [29]: we initialize $z_{T}$ with the noisy latent reference image.
Table 2 reports the comparison (see Ref normal-$\rightarrow$ Self). It could be found that our method surpasses the other state-of-the-art methods by a large margin (4.75 CLIP similarity and 1.23 FID). The superiority of CLIP similarity indicates that our method is able to better mimic the character of the background of another reference image, while the extremely low FID demonstrates that our method could generate new backgrounds consistently following the distribution of the advertising images.
Method Ref $\rightarrow$ Self Self $\rightarrow$ Self
Sim ($\uparrow$) FID ($\downarrow$) Sim ($\uparrow$) FID ($\downarrow$)
Blended Diffusion [1] 63.51 3.62 89.02 2.73
Paint-by-Example [26] 62.82 5.96 81.00 5.99
Stable ${}^{\dagger}$ [2] 57.10 7.92 76.10 7.29
${}^{\dagger}$ [29] 62.78 3.92 87.21 3.27
Ours 68.26 2.39 93.63 1.39
Table 2: Background generation conditioned on the reference image on BG-pair. Sim is short for the CLIP similarity (the generated image v.s. the reference image).
Figure 6: Qualitative Results. More visualizations are in the supplementary materials. Our method could generate backgrounds similar to the reference images in style, layout, elements, etc.
Moreover, we consider a simpler situation, where the product in the reference image is exactly the product requiring the background. Therefore, all the expected elements in the generated background could be spatially aligned with those in the reference image, so that the model does not need to understand the reference image deeply. As shown in Table 2 (see Self normal-$\rightarrow$ Self), our method still achieves the best performance in CLIP similarity and FID. Meanwhile the particularly outstanding metrics inspire us to improve the quality of generated background by selecting reference images containing similar-shaped products.
We further conduct a user study for a more focused and comprehensive comparison. We select five well-performed methods from above experiments, average the rank scores (five stands for the best and one for worst) and show the results in Table 3. It could be found that from the participant’s subjective perspective, almost every time the background generated by our method is picked as the most similar one. Moreover, we are encouraged that our method also ranks first in quality. This is more than challenging as pursuing the similarity in the layout and elements might hurt the harmony, especially when the shapes of the product exhibit great differences in the two images. Besides, even without PG, our method outperforms other counterparts in quality, which also proves the effect of the category commonality.
Method Similarity ($\uparrow$) Quality ($\uparrow$)
${}^{\dagger}$ [29] 2.39 2.61
Blended Diffusion [1] 2.53 2.77
Paint-by-Example [26] 2.03 2.35
[…]
Ours 4.88 3.84
[…]
We further visualize the qualitative comparison in Fig. 6. It could be found from the first row that our method could mimic the abstract style of the reference image and generate a similar background adaptively. Meanwhile when the background of the reference image is a concrete scene (the last two rows), our method is able to generate background similar in layout, elements, etc.
Category-Wise Generator Personality-Wise Generator Initial Image Ref $\rightarrow$ Self
Sim ($\uparrow$) FID ($\downarrow$)
(a) $\checkmark$ $\checkmark$ 66.17 4.01
(b) $\checkmark$ $\checkmark$ 67.50 2.76
(c) $\checkmark$ $\checkmark$ 66.51 2.73
(d) $\checkmark$ $\checkmark$ $\checkmark$ 68.26 2.39
Table 4: Ablation study on different modules of our method.
5.4 Ablation Study
Effectiveness of CG: To evaluate the effect of the category commonality on personalized background generation, we replace CG with the original ControlNet inpainting model. Then a significant performance deterioration is observed from line (a) and (d) in Table 4. It shows the necessity of the category commonality even with a reference image given, as it forms the basis of the generated background.
Effectiveness of PG: We remove PG and initialize $z_{T}$ with the latent reference image with noise $\epsilon\sim\mathcal{N}(0,1)$ added for $T$ times, to delve into the role of PG in extracting personalized information. From line (b) and (d) in Table 4, this results in the drop of CLIP similarity by $0.76\%$, and the increase of FID by 0.37. It is because the model fails to recognize and leverage the background information explicitly from the reference image.
Impact of Initial Latent Images: We try not initializing $z_{T}$ during inference to study the effect of initializing $z_{T}$ with the noisy reference image. As shown in line (c) and (d) in Table 4, this has a negative influence on the similarity and FID. Thus we conclude that a proper initialization could benefit the quality of generated images.
Comparison on Different Augmentation Strategies: As illustrated in the 3rd column in Fig. 7, when no perturbation strategy is adopted, an anomaly could be observed in where the product in the reference image is originally placed (see the red rectangle). This is alleviated when we employ mask dilation during training. However, as shown in the 4th column in Fig. 7, some elements still look like blunt copies (see the red rectangle). When the mixup strategy is further added, the generated background is finally similar but not identical to the reference background.
Figure 7: Comparison on different augmentation strategies.
6 Conclusion
In this work, we focus on several practical challenges in the e-commerce product background generation. Firstly we improve the efficiency of large-scale generation by a Category-Wise Generator, where we map the common style of each category into a unique identifier by the mask-guided cross attention layers. Furthermore we maintain the personalized styles for some specific brands from reference images effectively. We propose a Personality-Wise Generator and adopt a background perturbation strategy to prevent the copy-and-paste. Finally, we release the first large-scale product background generation dataset BG60k. Experiments show that our method could generate high-quality backgrounds for products from different categories, and backgrounds similar in personalized style to reference images if given.
Appendix A Detailed Analysis on BG60k
The distribution of the top 100 categories in BG60K is illustrated in Fig. 8. It could be found that the categories in BG60K follow the long-tail challenging distribution. It results from the strategy in which we construct BG60k. When collecting images for BG60k, we sort the products in descending order of impressions within one week for better image quality, and retain the top products to ensure the total impressions over 500 million. The high CTR advertising images are then obtained only from these products. Therefore, the final distribution of categories in BG60k follows the real long-tailed category distribution of images with high impressions on the E-commerce platform.
Appendix B Diverse Qualitative Results
In this section we show more diverse qualitative results. Firstly, given the same product, we visualize the generated images conditioned on the different reference images. As illustrated in Fig. 9, all of the generated images share the similar style to the corresponding reference images, which demonstrates the flexibility of our method.
Meanwhile, given the same reference image, we visualize the images generated for different products. As shown in Fig. 10, the generated backgrounds are consistent in layout and elements, but have slight difference in some details. It turns out that our method is able to generated proper backgrounds for different products adaptively, without lost of the personalized styles of the reference backgrounds.
Finally, we provide more qualitative comparisons on background generation conditioned on the reference image in Fig. 11 (see the next page). Our method shows great superiority in generating both artistically stylized backgrounds (the first two rows) and photorealistic backgrounds (the last four rows).
Figure 8: The distribution of the top 100 categories in BG60k. Due to the limited space, we denote each category as an temporary id. The mapping between the category name and id will be released with BG60k.
Figure 9: (a) Product image. (b) Different generated images conditioned on different reference images.
Figure 10: (a) Reference image. (b) Different generated images for different products.
Figure 11: More qualitative comparisons.